OpenStudy (anonymous):

3(5)+18 5=-6 only 5=6 only 5=-6 or 5=6 5=-18 or 5=18

OpenStudy (anonymous):

uh...no

OpenStudy (anonymous):

3*5 = 15 + 18 = 33

OpenStudy (anonymous):

i meant = 18 ?

OpenStudy (anonymous):

why are you moving 18 at all?

OpenStudy (anonymous):

3(5)=18

OpenStudy (anonymous):

is your questions 3(5) = 18 or 3(5) +18...?

OpenStudy (anonymous):

=18 the 1st one

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

yeah i messed up on the prob . its my bad

OpenStudy (anonymous):

The answer is false

OpenStudy (anonymous):

it doesn't work O_O

OpenStudy (anonymous):

\[15\neq18\]

OpenStudy (anonymous):

whaaa ? seriously ?

OpenStudy (anonymous):

dayumn . they gave me a problem to have no answer ?

OpenStudy (anonymous):

It happens :)

OpenStudy (anonymous):

Don't over think this.

OpenStudy (anonymous):

oh wow , okaay thaanks for letting me know , makes me feel alot better . :)

OpenStudy (anonymous):

I'll take a medal of ya though :P

OpenStudy (anonymous):

new to this i dont really know how. ? sorry if you tell me ill do it ?

OpenStudy (anonymous):

(meaning just click best response lol, won't get anything)

OpenStudy (anonymous):

best responce on any of your replies ?

OpenStudy (anonymous):

response**

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and after it's best to close the thread so other viewers can show theres.

OpenStudy (anonymous):

okaaay:) gottcha . thanks aloot :) hope to see you helping me out again sometime :)

OpenStudy (anonymous):

Of coarse :P

OpenStudy (anonymous):

Now I have to go help someone with their calculus problem O_O

OpenStudy (anonymous):

okay :) ttyl then